OpenStudy (anonymous):

Will give medal to best answer A 12-gallon gas tank is 80% full. Which proportion could be used to determine how many gallons of gasoline are in the tank?

OpenStudy (nurali):

Let x = the number of gallons in the tank Since there are 12 gallons the amount of gas left is (80/100)*12 Knowing that (80/100)*12 = x You can divide both sides of the equation by 12 and get x/12 = 80/100

OpenStudy (anonymous):

Thank you soooo much

OpenStudy (nurali):

Anytime.